OpenStudy (anonymous):

why does this converge or diverge? Integral: from 0 to infinity (e^(2x))/((e^x)+1)dx

OpenStudy (anonymous):

Determine whether the following improper integrals converge or diverge; justify your answers: ... u = x3 - 1, the integral is limt -> 1+ (2/3)(x3 - 1)1/2|t2 = (2/3)71/2 - 0. ... But the integral of e-x from 1 to infinity does converge; it is equal to limR ...

OpenStudy (anonymous):

I think I am following but the answer I was given says that the integral diverges...it does not mention why...so I don't think I follow